Jeffrey left 4/7 of his wife and 1/4 of the remainder to his granddaughter. The part of his estate that is not accounted for,given as a factor is
1
Expert's answer
2021-04-20T07:42:47-0400

The remainder is "1-\frac 47=\frac 37,"

granddaughter's part is "\frac 37\cdot \frac 14=\frac{3}{28},"

Jeffrey's part is "1-\frac 47-\frac{3}{28}=1-\frac{19}{28}=\frac{9}{28}."